Please help with the following questions: Find the slope of the line dtermined by the following equation. 7x - y = 63
2)Determine whether the line PQ is parallel or pendicular (or neither)to a line slope of -2.
P (3,4) Q (7,6)
3) Find the values. f(x) = x2 + 2
f(3) =
f(-3) =

Answer by Alan3354(69443)

Find the slope of the line dtermined by the following equation. 7x - y = 63
y = 7x - 63
Slope, m, = 7
---------------
2)Determine whether the line PQ is parallel or pendicular (or neither)to a line slope of -2.
P(3,4) Q(7,6)
Slope m = diffy/diffx
m = (6-4)/(7-3) = 2/4
m = 1/2
The slope of lines perpendicular is the negative inverse, = -2
So they're perpendicular.
------------------
3) Find the values. f(x) = x^2 + 2
f(3) = 3^2 + 2
= 11
-------
f(-3) = (-3)^2 + 2
= 11